OpenStudy (anonymous):

Solve |5x – 15| = 10 Answers x = {-5, 1} x = {-1, 5} x = {1, 5} No Solution

OpenStudy (anonymous):

I totally forgot how to do this ! xD

OpenStudy (anonymous):

-10 = 5x +15 = 10 subtract 15 from both 10's -25 = 5x = -5 -5 = x = -1 No solution since I'm not seeing -1 and -5